# Shane Cooper and Ilia Kamotski

### v1

@@ 2 Problem formulation
Assume that, for some $m\in\mathbb{N}$, $I=(b_{m},a_{m+1})$ is a gap of $\sigma_{0}$ Let $\varepsilon_{0}:=(a_{m+1}-b_{m})/2C_{b}$. Then, the set
@@ 4 Some properties of the reduced Maxwell operators $\mathcal{B}(\theta)$
$\forall\theta,\theta^{\prime}\in\square^{*},\forall v\in V(\theta),\quad\inf_{v^{\prime}\in V(\theta^{\prime})}\|v-v^{\prime}\|_{H^{1}(\square;\mathbb{C}^{2})}\leq L_{V}|\theta-\theta^{\prime}|\|v\|_{H^{1}(\square;\mathbb{C}^{2})}$
@@
(i). Suppose $\theta\in\square^{*}\backslash\{0\}$, then $v\in V(\theta)$ if, and only if,
@@
$\langle e^{-\mathrm{i}\theta\cdot y}\mathrm{div}\,{v}\rangle=0\quad\text{and}\quad\langle e^{-\mathrm{i}\theta\cdot y}\mathrm{curl}\,{v}\rangle=0.$ (4.5)
@@
$\Delta a=f_{1}\quad\text{and}\quad\Delta b=f_{2}\quad\text{in $\square$};$
@@
$\Delta r_{p}$ $=F_{p}:=(\chi_{R}-1)e^{\mathrm{i}(\theta^{\prime}-\theta)\cdot y}\Big{(}2\mathrm{i}(\theta^{\prime}-\theta)\cdot\nabla p-|\theta^{\prime}-\theta|^{2}p\Big{)}+\chi_{R}e^{\mathrm{i}\theta^{\prime}\cdot y}\frac{D_{p}}{|R|},$
@@
$\Delta p=(\mathrm{i}\theta^{\prime}\cdot c)e^{\mathrm{i}\theta^{\prime}\cdot y}\frac{\chi_{R}}{|R|}\quad\&\quad\Delta q=(\mathrm{i}\theta^{\prime\perp}\cdot c)e^{\mathrm{i}\theta^{\prime}\cdot y}\frac{\chi_{R}}{|R|}.$
@@
$\left\|\nabla p+e^{\mathrm{i}\theta^{\prime}\cdot y}(\mathrm{i}\theta^{\prime}\cdot c)\frac{\mathrm{i}\theta^{\prime}}{|\theta^{\prime}|^{2}}\right\|_{H^{1}(\square)}\leq\frac{C}{\sqrt{|R|}}|\theta^{\prime}||c|,\quad\&\quad\left\|\nabla^{\perp}q+e^{\mathrm{i}\theta^{\prime}\cdot y}(\mathrm{i}\theta^{\prime\perp}\cdot c)\frac{\mathrm{i}\theta^{\prime\perp}}{|\theta^{\prime}|^{2}}\right\|_{H^{1}(\square)}\leq\frac{C}{\sqrt{|R|}}|\theta^{\prime}||c|.$
@@
$b(v_{\theta},\tilde{v})+c(v_{\theta},\tilde{v})$ $=\langle F,\tilde{v}\rangle,\quad\forall\tilde{v}\in V(\theta);$ (4.8)
$b(v_{\theta^{\prime}},\tilde{v}^{\prime})+c(v_{\theta^{\prime}},\tilde{v}^{\prime})$ $=\langle F,\tilde{v}^{\prime}\rangle,\quad\forall\tilde{v}^{\prime}\in V(\theta^{\prime});$ (4.9)
@@
$Q(v_{\theta}-\mathcal{P}(\theta)v_{\theta^{\prime}},\tilde{v})=\langle F,(I-\mathcal{P}(\theta^{\prime})\tilde{v}\rangle,\quad\forall\tilde{v}\in V(\theta);$ (4.10)
Indeed, by (4.8), (4.9), and the identity $Q(\mathcal{P}(\theta)v_{\theta^{\prime}},\tilde{v})=Q(v_{\theta^{\prime}},\mathcal{P}(\theta^{\prime})\tilde{v})$ it follows that
$Q(v_{\theta}-\mathcal{P}(\theta)v_{\theta^{\prime}},\tilde{v})$ $=\langle F,\tilde{v}\rangle-Q(\mathcal{P}(\theta)v_{\theta^{\prime}},\tilde{v})=\langle F,\tilde{v}\rangle-Q(v_{\theta^{\prime}},\mathcal{P}(\theta^{\prime})\tilde{v})$
@@
$\|\big{(}I-\mathcal{P}(\theta^{\prime})\big{)}\tilde{v}\|_{+}\leq C|\theta-\theta|\|\tilde{v}\|_{+},\quad\forall\tilde{v}\in V(\theta),\forall\theta\in\square^{*},$ (4.12)
@@ 5 Photonic crystals which exhibit gaps in the Bloch spectrum on the critical light-line
$\sigma_{0}=\overline{\bigcup_{n=1}^{\infty}[a_{n},b_{n}]},\quad a_{n}:=\min_{\theta\in\square^{*}}\lambda_{n}(\theta),b_{n}:=\max_{\theta\in\square^{*}}\lambda_{n}(\theta).$
@@ 5.1 One-dimensional photonic crystal fibres
Let us suppose we restrict ourselves to one-dimensional photonic crystal fibres, that is our dielectric media has additional homogeneous axis, say the $x_{2}$-axis. That is, $R$ can be assumed to be of the form $[-\pi,-\pi+h)\times[-\pi,\pi)$, for some $0<h<\pi$; i.e. the dielectric permittivity $\epsilon$ (see (2.1)) on the periodic cell $\square$ is of the form
$\epsilon(x_{1},x_{2},x_{3})=\left\{\begin{matrix}\epsilon_{0},&x_{1}\in[-\pi,-\pi+h)\\[1.00006pt] 1,&x_{1}\in[-\pi+h,\pi).\end{matrix}\right.$
Let us rewrite the spectral problem on the critical line in this case (cf. (2.8) and (2.21)). Let, $\lambda=\omega^{2}$ be an eigenvalue of $\mathcal{B}(\theta)$, for some $\theta\in\square^{*}$, with eigenfunction $u\in V(\theta)$. Without loss of generality $\theta=(\Theta,0)$, $\Theta\in[-\frac{1}{2},\frac{1}{2})$ and $u$ depends only on $x_{1}$, i.e. $u(x_{1},x_{2})\equiv u(x_{1})$.777 In general, we are considering plane EM-waves of the form (see (2.2)) $\tilde{E}(x)=E(x_{1})e^{i(k_{2}x_{2}+\omega kx_{3}-\omega t)},\quad\text{and}\quad\tilde{H}(x)=H(x_{1})e^{i(k_{2}x_{2}+\omega kx_{3}-\omega t)},$ for $\theta$-quasi-periodic amplitudes, $E,H$, with $\theta\in[-\frac{1}{2},\frac{1}{2})$. However, we can suppose without loss of generality $k_{2}=0$; indeed, for the general wave number $(k_{2},\omega k)$, we can always rotate the one-dimensional PCF so that $(k_{2},\omega k)$ is an homogeneous axis of the media (that we would denote as the third axis). Thus, we are propagating down one homogeneous axis but not the other.. Notice that if $\phi\in V(\theta)$ and $\phi(x_{1},x_{2})\equiv\phi(x_{1})$ then $\phi$ is a constant vector for $x_{1}\in[h,\pi)$. Indeed, here $\mathrm{div}\,{\phi}=\phi_{1}^{\prime}$ and $\mathrm{curl}\,{\phi}=\phi_{2}^{\prime}$, then $\phi$ is constant in $(h,\pi+h)$ and
@@
$-u_{1}^{\prime\prime}=\lambda\gamma u_{1}\quad\text{in $(-\pi,-\pi+h)$},\quad\text{and}\quad-u_{2}^{\prime\prime}=\lambda\gamma u_{2}\quad\text{in $(-\pi,-\pi+h)$}.$
@@
$-u_{1}^{\prime\prime}=\Lambda u_{1}&\qquad\text{in $(l_{1},l_{2})$},\\ e^{-\mathrm{i}\pi\Theta}u_{1}^{\prime}(l_{2})-u_{1}^{\prime}(l_{1})&=(2\pi-h)\Lambda u_{1}(l_{1}),\\ e^{-\mathrm{i}\pi\Theta}u_{1}(l_{2})&=u_{1}(l_{1});\end{aligned}\qquad\right\}$ (TE)
@@
$-u_{2}^{\prime\prime}=\Lambda u_{2}&\qquad\text{in $(l_{1},l_{2})$},\\ e^{-\mathrm{i}\pi\Theta}u_{2}^{\prime}(l_{2})-u_{2}^{\prime}(l_{1})&=\frac{(2\pi-h)}{1+\gamma}\Lambda u_{2}(l_{1}),\\ e^{-\mathrm{i}\pi\Theta}u_{2}(l_{2})&=u_{2}(l_{1}).\end{aligned}\qquad\right\}$ (TM)
@@
Notice that (TE) and (TM) are both problems of the same type; this type of problem also appears as the asymototic limit of the one-dimensional high-contrast double porosity model studied by one of the authors in [2] and [3]. In particular, it was shown in [3, Theorem 3.1] that the spectral band functions $E_{n}(\Theta)$ of the associated operator has the usual Floquet-type properties; namely:
@@
  • For $n$ odd, $E_{n}$ is strictly increasing in $\Theta$ from 00 to $\frac{1}{2}$;
  • For $n$ even, $E_{n}$ is strictly decreasing in $\Theta$ from 00 to $\frac{1}{2}$;
@@
    where, for $j\in\{\mathrm{TE},\mathrm{TM}\}$, $\mathrm{Ran}E_{n}^{(j)}$, the range of $E_{n}^{(j)}$, equals to $[E_{n}^{(j)}(0),E_{n}^{(j)}(\frac{1}{2})]$ for $n$ odd and equals to $[E_{n}^{(j)}(\frac{1}{2}),E_{n}^{(j)}(0)]$ for $n$ even.
@@
$\Lambda_{n}^{(j)}(\tfrac{1}{2})<\Lambda^{(j)}_{n+1}(\tfrac{1}{2})\quad\text{for some odd $n$},$
@@
$\Lambda^{(j)}_{n}(0)<\Lambda^{(j)}_{n+1}(0)\quad\text{for some even $n$}.$
@@ 5.2 ARROW fibres: two-dimensional photonic crystals with inclusions have vanishing volume faction
$-\Delta p=\chi_{R}f\quad\text{in $\square$.}$
Here $\chi_{R}:L^{2}(R)\rightarrow L^{2}(\mathbb{R}^{2})$ is the operator that extends the input function by zero into $\mathbb{R}^{2}\backslash R$. From (4.1), it follows that $V(\theta)$ has the following representation in terms of $A_{\theta}$:
@@
$-\Delta p=\chi_{\Omega}f\quad\text{in $\mathbb{R}^{2}$},\quad p(x)=-\frac{d}{2\pi}\ln{|x|}+o(1)\ \text{as $|x|\rightarrow+\infty$},$
@@
$\mathfrak{b}(f,\tilde{f}):=\gamma^{-1}\int_{\Omega}f\cdot\overline{\tilde{f}}-\int_{\Omega}\left[\Big{(}(Af_{1})_{,22}+(Af_{2})_{,12}\Big{)}\overline{\tilde{f}_{1}}+\Big{(}(Af_{1})_{,21}+(Af_{2})_{,11}\Big{)}\overline{\tilde{f}_{2}}\right],$
@@
$-\Delta G=\delta\quad\text{in $\square$},\qquad\text{for $\delta$ the Dirac delta distribution.}$.
@@
As such we have a problem of the type studied, by the authors, in [5, Chapter 3] and we shall follow the spirit the arguments used therein; however, for this particular, example we shall prove slightly more.
@@
$\mathfrak{b}[f]=\gamma^{-1}\int_{\Omega}|f|^{2}+\int_{\Omega}\big{(}|(Af_{1})_{,12}+(Af_{2})_{,11}|^{2}+|(Af_{1})_{,22}+(Af_{2})_{,12}|^{2}\big{)},\quad\forall f\in L^{2}(\Omega;\mathbb{C}^{2}).$
@@
$\mathfrak{b}(Nc,\tilde{g})=-\mathfrak{b}(c,\tilde{g}),\quad\forall\tilde{g}\in L^{2}_{0}(\Omega;\mathbb{C}^{2}).$
@@
$\mathfrak{b}(\tilde{g},(I+N)\tilde{c})=0,\quad\forall\tilde{g}\in L^{2}_{0}(\Omega;\mathbb{C}^{2}),\quad\forall\tilde{c}\in\mathbb{C}^{2}.$
@@
$\mathfrak{c}_{1}[(I+N)c+g]=\mathfrak{c}_{1}[c],\quad\forall\tilde{c}\in\mathbb{C}^{2}.$
@@
$\mu^{(\delta)}_{n}=\Big{(}1+\mathcal{O}\left(\ln(\pi/\delta)^{-1/2}\right)\Big{)}\Lambda^{(\delta)}_{n}(\theta)\qquad\text{as $\delta\rightarrow 0$,}$ (5.11)
@@
$\mathfrak{c}_{2}[(I+N)c]+\mathfrak{b}[(I+N)c]\leq\kappa_{1}^{2}\mathfrak{c}_{1}[c],\quad\forall c\in\mathbb{C}^{2}.$ (5.12)
@@
$\kappa_{3}\mathfrak{c}_{2}[f]\leq\mathfrak{b}[f],\quad\forall f\in L^{2}(\Omega;\mathbb{C}^{2}),$
@@
$\Lambda^{(\delta)}_{n}(\theta)$ $=\nu_{\theta,\delta}^{-1}\Lambda_{n},\quad n=1,2;$ (5.14)
$\Lambda^{(\delta)}_{n}(\theta)$ $=\Lambda_{n}:=\min_{\begin{subarray}{c}V\subset L^{2}_{0}(\Omega;\mathbb{C}^{2}),\\ \dim{V}=(n-2)\end{subarray}}\max_{g\in V}\frac{\mathfrak{b}[g]}{\mathfrak{c}_{2}[g]},\quad n=3,4,\ldots.$ (5.15)
@@
$\lambda^{(\delta)}_{n}(\theta)=\frac{1}{\delta^{2}\big{(}g_{\theta}(0)-(2\pi)^{-1}\ln{\delta}\big{)}}\Lambda_{n}+\mathcal{O}(\delta^{-2}|\ln{\delta}|^{-3/2}),\quad n=1,2,$
@@
$\lambda^{(\delta)}_{n}(\theta)=\delta^{-2}\Lambda_{n}+\mathcal{O}_{n}(\delta^{-2}|\ln{\delta}|^{-1/2}),\quad n\geq 3,$
@@
$\ΩA1=\left\{\begin{array}[]{lcl}-\frac{|x|^{2}}{4}+\frac{1}{4}&&\text{in $\Omega$}\\[1.99997pt] -\frac{1}{2}\ln{|x|}&&\text{in $\mathbb{R}^{2}\backslash\Omega$}\end{array}\right..$
@@
$\mathfrak{b}(\tilde{c},\tilde{g})=0,\quad\forall\tilde{c}\in\mathbb{C}^{2},\forall\tilde{g}\in L^{2}_{0}(\Omega;\mathbb{C}^{2}).$

### v2

@@ 2 Problem formulation
Assume that, for some $m\in\mathbb{N}$, $I=(b_{m},a_{m+1})$ is a gap of $\sigma_{0}$ . Let $\varepsilon_{0}:=(a_{m+1}-b_{m})/2C_{b}$. Then, the set
@@ 4 Some properties of the reduced Maxwell operators $\mathcal{B}(\theta)$
$\forall\theta,\theta^{\prime}\in\square^{*},\forall v\in V(\theta),\qquad\inf_{v^{\prime}\in V(\theta^{\prime})}\|v-v^{\prime}\|_{H^{1}(\square;\mathbb{C}^{2})}\leq L_{V}|\theta-\theta^{\prime}|\|v\|_{H^{1}(\square;\mathbb{C}^{2})}$
@@
[…]
@@
$\langle e^{-\mathrm{i}\theta\cdot y}\mathrm{div}\,{v}\rangle=0\qquad\text{and}\qquad\langle e^{-\mathrm{i}\theta\cdot y}\mathrm{curl}\,{v}\rangle=0.$ (4.5)
@@
$\Delta a=f_{1}\qquad\text{and}\qquad\Delta b=f_{2}\qquad\text{in $\square$};$
@@
$\Delta r_{p}$ $=F_{p}=(\chi_{R}-1)e^{\mathrm{i}(\theta^{\prime}-\theta)\cdot y}\Big(2\mathrm{i}(\theta^{\prime}-\theta)\cdot\nabla p-|\theta^{\prime}-\theta|^{2}p\Big)+\chi_{R}e^{\mathrm{i}\theta^{\prime}\cdot y}\frac{D_{p}}{|R|},$
@@
$\Delta p=(\mathrm{i}\theta^{\prime}\cdot c)e^{\mathrm{i}\theta^{\prime}\cdot y}\frac{\chi_{R}}{|R|}\qquad\&\qquad\Delta q=(\mathrm{i}\theta^{\prime\perp}\cdot c)e^{\mathrm{i}\theta^{\prime}\cdot y}\frac{\chi_{R}}{|R|}.$
@@
$\left\|\nabla p+e^{\mathrm{i}\theta^{\prime}\cdot y}(\mathrm{i}\theta^{\prime}\cdot c)\frac{\mathrm{i}\theta^{\prime}}{|\theta^{\prime}|^{2}}\right\|_{H^{1}(\square)}\leq\frac{C}{\sqrt{|R|}}|\theta^{\prime}||c|,\qquad\&\qquad\left\|\nabla^{\perp}q+e^{\mathrm{i}\theta^{\prime}\cdot y}(\mathrm{i}\theta^{\prime\perp}\cdot c)\frac{\mathrm{i}\theta^{\prime\perp}}{|\theta^{\prime}|^{2}}\right\|_{H^{1}(\square)}\leq\frac{C}{\sqrt{|R|}}|\theta^{\prime}||c|.$
@@
[…]
@@
$Q(v_{\theta}-\mathcal{P}(\theta)v_{\theta^{\prime}},\tilde{v})=\langle F,(I-\mathcal{P}(\theta^{\prime})\tilde{v}\rangle,\qquad\forall\tilde{v}\in V(\theta);$ (4.10)
[…]
@@
$\|\big(I-\mathcal{P}(\theta^{\prime})\big)\tilde{v}\|_{+}\leq C|\theta-\theta|\|\tilde{v}\|_{+},\qquad\forall\tilde{v}\in V(\theta),\forall\theta\in\square^{*},$ (4.12)
@@ 5 Photonic crystals which exhibit gaps in the Bloch spectrum on the critical light-line
$\sigma_{0}=\overline{\bigcup_{n=1}^{\infty}[a_{n},b_{n}]},\qquad a_{n}:=\min_{\theta\in\square^{*}}\lambda_{n}(\theta),b_{n}:=\max_{\theta\in\square^{*}}\lambda_{n}(\theta).$
@@ 5.1 One-dimensional photonic crystal fibres
Let us suppose we restrict ourselves to one-dimensional photonic crystal fibres, that is our dielectric media has additional homogeneous axis, say the $x_{2}$-axis. That is, $R$ can be assumed to be of the form $\pi,-\pi+h)\times[-\pi,\pi)$, for some $0<h<\pi$; i.e. the dielectric permittivity $\epsilon$ (see (2.1)) on the periodic cell $\square$ is of the form
$\epsilon(x_{1},x_{2},x_{3})=\left\{\begin{matrix}\epsilon_{0},&x_{1}\in[-\pi,-\pi+h)\\[0.85005pt] 1,&x_{1}\in[-\pi+h,\pi).\end{matrix}\right$.
Let us rewrite the spectral problem on the critical line in this case (cf. (2.8) and (2.21)). Let, $\lambda=\omega^{2}$ be an eigenvalue of $\mathcal{B}(\theta)$, for some $\theta\in\square^{*}$, with eigenfunction $u\in V(\theta)$. Without loss of generality $\theta=(\Theta,0)$, $\Theta\in[-\frac{1}{2},\frac{1}{2})$ and $u$ depends only on $x_{1}$, i.e. $u(x_{1},x_{2})\equiv u(x_{1})$. [^7] In general, we are considering plane EM-waves of the form (see (2.2)) $\tilde{E}(x)=E(x_{1})e^{i(k_{2}x_{2}+\omega kx_{3}-\omega t)},\qquad\text{and}\qquad\tilde{H}(x)=H(x_{1})e^{i(k_{2}x_{2}+\omega kx_{3}-\omega t)},$ for $\theta$-quasi-periodic amplitudes, $E,H$, with $\theta\in[-\frac{1}{2},\frac{1}{2})$. However, we can suppose without loss of generality $k_{2}=0$; indeed, for the general wave number $(k_{2},\omega k)$, we can always rotate the one-dimensional PCF so that $(k_{2},\omega k)$ is an homogeneous axis of the media (that we would denote as the third axis). Thus, we are propagating down one homogeneous axis but not the other.. Notice that if $\phi\in V(\theta)$ and $\phi(x_{1},x_{2})\equiv\phi(x_{1})$ then $\phi$ is a constant vector for $x_{1}\in[h,\pi)$. Indeed, here $\mathrm{div}\,{\phi}=\phi_{1}^{\prime}$ and $\mathrm{curl}\,{\phi}=\phi_{2}^{\prime}$, then $\phi$ is constant in $(h,\pi+h)$ and
@@
$u_{1}^{\prime\prime}=\lambda\gamma u_{1}\qquad\text{in $(-\pi,-\pi+h)$},\qquad\text{and}\qquad-u_{2}^{\prime\prime}=\lambda\gamma u_{2}\qquad\text{in $(-\pi,-\pi+h)$}.$
@@
$\left.\begin{aligned} -u_{1}^{\prime\prime}=\Lambda u_{1}&\hskip 17.00024pt\text{in $(l_{1},l_{2})$},\\ e^{-\mathrm{i}\pi\Theta}u_{1}^{\prime}(l_{2})-u_{1}^{\prime}(l_{1})&=(2\pi-h)\Lambda u_{1}(l_{1}),\\ e^{-\mathrm{i}\pi\Theta}u_{1}(l_{2})&=u_{1}(l_{1});\end{aligned}\hskip 17.00024pt\right\}$ (TE)
@@
$\left.\begin{aligned} -u_{2}^{\prime\prime}=\Lambda u_{2}&\hskip 17.00024pt\text{in $(l_{1},l_{2})$},\\ e^{-\mathrm{i}\pi\Theta}u_{2}^{\prime}(l_{2})-u_{2}^{\prime}(l_{1})&=\frac{(2\pi-h)}{1+\gamma}\Lambda u_{2}(l_{1}),\\ e^{-\mathrm{i}\pi\Theta}u_{2}(l_{2})&=u_{2}(l_{1}).\end{aligned}\hskip 17.00024pt\right\}$ (TM)
@@
Notice that (TE) and (TM) are both problems of the same type; this type of problem also appears as the asymototic limit of the one-dimensional high-contrast double porosity model studied by one of the authors in [5] and [6]. In particular, it was shown in [6, Theorem 3.1] that the spectral band functions $E_{n}(\Theta)$ of the associated operator has the usual Floquet-type properties; namely:
@@
  • For $n$ odd, $E_{n}$ is strictly increasing in $\Theta$ from 0 to $\frac{1}{2}$;
  • For $n$ even, $E_{n}$ is strictly decreasing in $\Theta$ from 0 to $\frac{1}{2}$;
@@
    where, for $j\in\{\mathrm{TE},\mathrm{TM}\}$, $\mathrm{Ran}E_{n}^{(j)}$, the range of $E_{n}^{(j)}$, equals to $E_{n}^{(j)}(0),E_{n}^{(j)}(\frac{1}{2})]$ for $n$ odd and equals to $E_{n}^{(j)}(\frac{1}{2}),E_{n}^{(j)}(0)]$ for $n$ even.
@@
$\Lambda_{n}^{(j)}(\tfrac{1}{2})<\Lambda^{(j)}_{n+1}(\tfrac{1}{2})\qquad\text{for some odd $n$},$
@@
$\Lambda^{(j)}_{n}(0)<\Lambda^{(j)}_{n+1}(0)\qquad\text{for some even $n$}.$
@@ 5.2 ARROW fibres: two-dimensional photonic crystals with inclusions have vanishing volume faction
$\Delta p=\chi_{R}f\qquad\text{in $\square$.}$
Here $\chi_{R}:L^{2}(R)\rightarrow L^{2}(\mathbb{R}^{2})$ is the operator that extends the input function by zero into $\R\mathbb{R}^{2}\backslash$ R. From (4.1), it follows that $V(\theta)$ has the following representation in terms of $A_{\theta}$:
@@
$\Delta p=\chi_{\Omega}f\qquad\text{in $\mathbb{R}^{2}$},\qquad p(x)=-\frac{d}{2\pi}\ln{|x|}+o(1)\ \text{as $|x|\rightarrow+\infty$},$
@@
$\mathfrak{b}(f,\tilde{f})=\gamma^{-1}\int_{\Omega}f\cdot\overline{\tilde{f}}-\int_{\Omega}\left[\Big((Af_{1})_{,22}+(Af_{2})_{,12}\Big)\overline{\tilde{f}_{1}}+\Big((Af_{1})_{,21}+(Af_{2})_{,11}\Big)\overline{\tilde{f}_{2}}\right],$
@@
$\Delta G=\delta\qquad\text{in $\square$},\hskip 17.00024pt\text{for $\delta$ the Dirac delta distribution.}$
@@
As such we have a problem of the type studied, by the authors, in [8, Chapter 3] and we shall follow the spirit the arguments used therein; however, for this particular, example we shall prove slightly more.
@@
$\mathfrak{b}[f]=\gamma^{-1}\int_{\Omega}|f|^{2}+\int_{\Omega}\big(|(Af_{1})_{,12}+(Af_{2})_{,11}|^{2}+|(Af_{1})_{,22}+(Af_{2})_{,12}|^{2}\big),\qquad\forall f\in L^{2}(\Omega;\mathbb{C}^{2}).$
@@
$\mathfrak{b}(Nc,\tilde{g})=-\mathfrak{b}(c,\tilde{g}),\qquad\forall\tilde{g}\in L^{2}_{0}(\Omega;\mathbb{C}^{2}).$
@@
$\mathfrak{b}(\tilde{g},(I+N)\tilde{c})=0,\qquad\forall\tilde{g}\in L^{2}_{0}(\Omega;\mathbb{C}^{2}),\qquad\forall\tilde{c}\in\mathbb{C}^{2}.$
@@
$\mathfrak{c}_{1}[(I+N)c+g]=\mathfrak{c}_{1}[c],\qquad\forall\tilde{c}\in\mathbb{C}^{2}.$
@@
$\mu^{(\delta)}_{n}=\Big(1+\mathcal{O}\left(\ln(\pi/\delta)^{-1/2}\right)\Big)\Lambda^{(\delta)}_{n}(\theta)\hskip 17.00024pt\text{as $\delta\rightarrow 0$,}$ (5.11)
@@
$\mathfrak{c}_{2}[(I+N)c]+\mathfrak{b}[(I+N)c]\leq\kappa_{1}^{2}\mathfrak{c}_{1}[c],\qquad\forall c\in\mathbb{C}^{2}.$ (5.12)
@@
$\kappa_{3}\mathfrak{c}_{2}[f]\leq\mathfrak{b}[f],\qquad\forall f\in L^{2}(\Omega;\mathbb{C}^{2}),$
@@
$\Lambda^{(\delta)}_{n}(\theta)$ $=\nu_{\theta,\delta}^{-1}\Lambda_{n},\qquad n=1,2;$ (5.14)
$\Lambda^{(\delta)}_{n}(\theta)$ $=\Lambda_{n}:=\min_{\begin{subarray}{c}V\subset L^{2}_{0}(\Omega;\mathbb{C}^{2}),\\ \dim{V}=(n-2)\end{subarray}}\max_{g\in V}\frac{\mathfrak{b}[g]}{\mathfrak{c}_{2}[g]},\qquad n=3,4,\ldots.$ (5.15)
@@
$\lambda^{(\delta)}_{n}(\theta)=\frac{1}{\delta^{2}\big(g_{\theta}(0)-(2\pi)^{-1}\ln{\delta}\big)}\Lambda_{n}+\mathcal{O}(\delta^{-2}|\ln{\delta}|^{-3/2}),\qquad n=1,2,$
@@
$\lambda^{(\delta)}_{n}(\theta)=\delta^{-2}\Lambda_{n}+\mathcal{O}_{n}(\delta^{-2}|\ln{\delta}|^{-1/2}),\qquad n\geq 3,$
@@
$A1=\left\{\begin{array}[]{lcl}-\frac{|x|^{2}}{4}+\frac{1}{4}&&\text{in $\Omega$}\\[1.7pt] -\frac{1}{2}\ln{|x|}&&\text{in $\mathbb{R}^{2}\backslash\Omega$}\end{array}\right.$.
@@
$\mathfrak{b}(\tilde{c},\tilde{g})=0,\qquad\forall\tilde{c}\in\mathbb{C}^{2},\forall\tilde{g}\in L^{2}_{0}(\Omega;\mathbb{C}^{2}).$